OpenStudy (anonymous):

At the beginning of the first day, Ann has $25$ dollars. By the end of the first day, Ann has $\frac{2}{3}$ of the money that she had at the beginning of the first day. By the end of the second day, Ann has $\frac{4}{7}$ of the money that she had at the end of the first day. How much money does Ann have at the end of the second day? Express your answer (in dollars) as a fraction.

OpenStudy (help!!!!):

\[\frac{2}{3}\] \[\frac{4}{7}\]

OpenStudy (help!!!!):

So the first day is $25*2/3=first day

OpenStudy (anonymous):

ok

OpenStudy (help!!!!):

after you find the first day multiply it by 4/7

OpenStudy (help!!!!):

So basically you keep day one in a fraction DAY 1: $25*2/3=50/3 DAY 2:50/3*4/7=200/21 which simplifies to 9 11/21

OpenStudy (anonymous):

oh ok